the sum of three peoples age is 57 the oldest is three times as old as the youngest the middle person is 15 years younger than the oldest what are the ages of all three kids
Answer by jorel555(1290)

Let n be the youngest. Then:
n+3n+(3n-15)=57
7n=72
n=72/7
3n=216/7
(3n-15)=111/7
The youngest is 72/7 years old, the middle is 111/7 years old, and the oldest is 216/7 years old!!!!!!!!!!